# Privacy-Preserving Edge Federated Learning for Intelligent Mobile-Health Systems

### v1

@@ Abstract
Index Terms:
Edge Federated Learning, Mobile-Health Technologies, Privacy-Preserving Machine Learning
I Introduction
Over the past few decades, we have been witnessing a dramatic increase in healthcare costs. In the United States of America, for instance, the healthcare costs have risen from 27.2 billion dollars (only 5 percent of GDP) in 1960 to 41.2 billion dollars (19.7 percent of GDP) in 2020 [1, 2]. Current practice in the healthcare system places a heavy burden on the society, in terms of healthcare expenditures, due to its treatment-oriented philosophy. Mobile-health and wearable technologies offer a promising solution to pervasive healthcare by removing the constraints with respect to time and location [3].
@@ I Introduction
The majority of the state-of-the-art detection techniques for such adverse health events, i.e., epileptic seizures, strokes, heart attacks, adopt modern ML algorithms, e.g., deep neural networks [10, 11, 12, 13]. However, training reliable deep neural network models requires a large amount of data, i.e., the so-called Big Data. As a result, the training is often performed based on the data collected from many patients, which are stored and distributed on patients’ mobile devices and wearable sensors. Traditionally, all training data samples are assumed to be present where the ML algorithm runs, i.e., the assumption was that the learning algorithm has access to all training data samples throughout the entire learning process. Therefore, to learn a model by adopting conventional solutions, each party holding a portion of the data is required to transfer all its training data in raw format to a center for the learning processing. However, such an approach raises major privacy concerns in connection to sharing sensitive personal/patient data [14, 15].
@@
The paper is structured as follows. Section II reviews the state of the art in FL, including in the context of the IoT systems and health applications, with a main focus on privacy. Section III illustrates the proposed framework for privacy-preserving edge FL, designed for resource-constrained mobile-health and wearable technologies. Section IV evaluates the proposed framework on real-world IoT applications, including implementation of our proposed framework on Amazon’s AWS cloud. Finally, Section V serves as the conclusion of this article.
II State of the Art
@@
In the past few years, a large number of studies consider FL to address their concerns related to privacy and overhead, particularly in the healthcare domain, [39, 40, 12]. FL methods are usually based on neural network algorithms, but the term has also been used for other ML algorithms recently, e.g., for tree-based algorithms [41, 42], for algorithms that are both gradient-based and tree-based [43, 44, 45, 46, 47], or for other algorithms [48, 49]. Several literature reviews of the state-of-the-art FL techniques are available in [50, 51, 52].
@@
A comprehensive study on the security and privacy of FL is performed in [86]. The authors extensively discuss the vulnerabilities and threats in FL. For instance, one concern in FL is data poisoning [87], which is relevant when the attacker incorporates malicious data points to training data in order to maximize the error. In another example, the global model can be used to infer details about training data. Membership inference attack [80, 81] may be used to determine if a certain record has existed among the training data samples.
@@
III Privacy-Preserving FL
In this section, we describe our proposed privacy-preserving FL framework for mobile-health and wearable technologies. Section III-A will briefly discuss the privacy model we consider in this paper. Then, in Section III-B, we discuss our proposed privacy-preserving edge FL framework for mobile-health and wearable technologies.
III-A Privacy Model
@@
III-B Privacy-Preserving Edge FL
We shall now present our proposed privacy-preserving edge FL in two steps. In the first step, in Section III-B1, we illustrate the initialization procedure among healthcare centers to train an initial base neural network model in a privacy-preserving fashion. Then, in Section III-B2, we discuss our privacy-preserving FL framework for mobile-health and wearable technologies, which reduces the complexity by exploiting the the initial base neural network model trained in Section III-B1.
III-B1 FL Base Model
In the initialization phase, we assume $N$ hospitals or healthcare centers are involved in training an initial generic model for epileptic seizure detection. The training data is distributed among all healthcare centers involved and the training is required to be performed in a secure and distributed fashion, since the personal/patient medical data is often considered sensitive and may not leave hospital premises due to regulatory restrictions. This scenario is shown in Figure 1.
@@ III-B1 FL Base Model
Given a deep neural network model and the data, the goal is to train a robust and reliable model based on the training data available on all healthcare centers. The model/hypothesis is shown by $h_{\theta}(X)$, in which $\theta$ is the set of the network’s parameters that are optimized during the process of learning, and $X$ is the input to the network. The objective is to minimize the loss represented by the following function:
@@
where $m$ represents the number of training data samples. $X^{(i)}$ is the vector of attributes for the $i^{\textrm{th}}$ training data sample, and $y^{(i)}$ is the target value for the $i^{\textrm{th}}$ training data sample. Having the loss function $J(\theta)$, the model/hypothesis parameters $\theta$, the training data $X$, and the labels $y$, we compute the gradients for updating the model/hypothesis parameters. Having the gradients, we update the model parameters based on an optimization algorithm, e.g., the Adam optimizer. This process summarizes the conventional centralized training procedure for DNNs.
@@
The standard FL scheme, however, while reducing the privacy concerns, does not guarantee privacy. Therefore, the data-holder parties may not directly share their contribution, partial gradients, with other parties for privacy and legal concerns [76, 77, 78, 79]. Several studies incorporate noise to data-holder parties’ contribution in order to provide differential privacy [84]. However, this degrades the prediction performance as there is a trade-off between privacy and data utility in differential privacy [85]. In contrast, here, the optimizer merely requires the global gradients to update the network parameters and we do not require parties to share their partial gradients in their raw format. Therefore, in this study, for calculating the global gradient, we employ a secure aggregation technique. In essence, our proposed framework masks the contributions of each data-holder party to preserve privacy. Towards this, each party includes its local masks to the partial gradients and only then shares the results with the mediator (a central server orchestrating the training process):
@@ III-B2 FL on Edge
In the first step, similar to the previous section, the key exchange protocol via the Diffie-Hellman algorithms [108] is adopted to exchange secret shares among the parties in a pair-wise fashion. The Diffie-Hellman key exchange scheme is, however, relatively complex when it comes to mobile-health and wearable technologies. To alleviate this, instead of executing the Diffie-Hellman key exchange scheme in every iteration of the learning process, we propose to share a secret random seed pair-wise among the parties in the initialization phase. This seed will then be used to generate secret masks during the entire process of privacy-preserving FL.
@@
The key challenge here is that training this model from scratch incurs major overheads for resource-constrained mobile-health and wearable technologies. On the one hand, while the initial generic neural network model trained collaboratively by the healthcare centers (in Section III-B1) is tailored to the high-quality data from the high-tech hospital equipment, this model can still be used as a basis for the data acquired by the mobile-health and wearable technologies. This will allow faster convergence of the model, hence reducing the communication and computation overheads of the training.
Unfortunately, training the entire neural network, even starting from the already trained models in the initial step, is still computationally infeasible for the resource-constrained mobile-health and wearable technologies. To address this challenge, inspired by the transfer learning techniques, we propose to only refine the model partially, where only the last few layers are fine-tuned to the data acquired by the mobile-health and wearable devices. More specifically, the initial generic model, denoted by $M$, is already available. As shown in Figure 2, the model $M=[B\rightarrow H]$ consists of two parts: the base architecture, denoted by $B$, and the head architecture, denoted by $H$. To reduce the overheads of training on resource-constrained mobile-health and wearable devices, we only train the head architecture $H$ locally on the mobile-health devices in a federated and privacy-preserving fashion. The base architecture $B$, while trained on high-quality data from high-tech hospital equipment, is still able to capture patterns which are also relevant in the case of the data acquired by the mobile-health and wearable devices. Mathematically, this is captured by the following equation:
@@
Finally, our proposed framework also provides the possibility of personalization, to enable the realization of the precision medicine paradigm. Once the FL process discussed above is completed, the head $H$ can be immediately refined based on the data locally available to each patient. In this way, the parameters in the head $H$ architecture will be fine-tuned to the specific profile of each patient, enabling precision detection of health pathology, as we show experimentally in the next section.
In summary, our proposed privacy-preserving edge FL framework is specifically designed for resource-constrained IoT, mobile-health, and wearable technologies, not only from the FL perspective, but also from a security perspective. In the next session, we will evaluate our proposed privacy-preserving edge FL framework with respect to prediction performance, privacy, and overheads.
IV Experimental Setup and Results
This section presents the experimental results and evaluation of our proposed framework with respect to classification performance, overhead, and privacy, which are the main criteria for the assessment of privacy-preserving data mining algorithms [110]. In Section IV-A, we discuss the experimental setup that includes the specification of the data in our experiments and the architecture of our neural network. In Section IV-B, we evaluate our proposed privacy-preserving edge FL framework in terms of overheads, complexity, latency, and privacy, based on an implementation of the proposed framework on Amazon’s AWS cloud platform.
IV-A Experimental Setup
IV-A1 Epilepsy Dataset
In this work, we consider the CHB-MIT database [21, 22] that contains Electroencephalogram (EEG) signals from 23 epilepsy patients with intractable seizures. All recordings are collected from children and young adults in the 1.5–22 age range. For collecting data, the international 10-20 system is used, which exploits EEG electrodes with a frequency of 256 Hz and 16-bit resolution. The dataset is annotated by the medical experts and contains a total of 182 seizures. These EEG signals are sampled at $F_{s}=256$ Hz, with 16-bit resolution. Here, we consider only the two channels $T7F7$ and $T8F8$ in the e-Glass wearable system [7], which have been shown to be important for the detection of epileptic seizures.
IV-A2 Seizure Detection Performance Metrics
@@
where $FP$, $TN$, $TP$, and $FN$ capture the number of false-positive samples, true-negative samples, true-positive samples, and false-negative samples, respectively.
IV-A3 Deep Neural Network’s Architecture
As described in Section III, we have two phases of FL. In the first phase, since the training is performed on decentralized data stored in several hospitals, the assumption is that we do not have limitations in regard to the resources required for training our model. Therefore, we can have more trainable parameters in the network’s base if increasing the size of the base improves the classification performance. On the other hand, in the second phase, since the training is performed on resource-constrained devices, we desire to have a lower number of network parameters to train. To this end, we designed our network to have three convolution layers with 135,052 trainable parameters in the base to extract the features from input EEG signals better. For the head of our network, we merely employed one dense layer with 8,234 trainable parameters.
@@ IV-A3 Deep Neural Network’s Architecture
Figure 3 shows the architecture of the base and head of the neural network we used in our experiments. [^1] The network architecture was plotted using [111]. We designed the base to have three layers of one-dimensional convolution, each followed by a batch normalization, an activation function, and a dropout. The convolution kernel size and stride are set to 2, and the number of filters for the first, second, and third layers are set to 512, 128, and 4, respectively. The activation functions in the base are rectified linear unit (ReLu) functions, and the dropout rates are set to 30%. For the head of the network, we consider a dense layer with eight units, followed by batch normalization, ReLu activation function, and a dropout. Then, we have an output layer that contains two units with softmax activation functions.
TABLE I: Quantitative comparison of the proposed framework against the baseline techniques (average of 100 runs)
@@
IV-B Experimental Results
IV-B1 Prediction Performance
In this section, we evaluate our proposed edge FL framework in terms of prediction performance and for the case of epileptic seizure detection. We perform the experiments 100 times and report the average value in Table I. Consistent with the CHB-MIT dataset, we consider 23 patients and each with his/her own data locally and the e-Glass system [7]. We consider four baseline approaches as follows:
@@ IV-B1 Prediction Performance
Next, our proposed framework also provides the possibility of personalization. In this setting, we follow the same exact procedure as discussed in our proposed framework. After the FL is finalized, we fine-tune the head architecture with the local data of the patient to develop a personalized model for that patient. We have repeated this experiment for all patients to obtain a personalized model for each patient. The last row of Table I shows the average results of the personalization for all patients. The results indicate that the personalized models outperform their generic counterpart by $7.0\%$ and $6.2\%$ in terms of accuracy and F1-score, respectively. This demonstrates that our proposed framework supports personalization to significantly improve the prediction performance.
Finally, here, we show that the proposed approach is not limited to the architecture considered in Figure 3. To demonstrate this, we consider two state-of-the-art architectures, namely Transformer [112] and ResNet [113]. Our Transformer model consists of one encoder block, including multi-head attention with eight heads and an embedding size of 32. The feedforward part consists of two 1-dimensional convolution layers with kernel sizes of 1, with 8 and 16 filters and strides of 2. Each layer consists of 1-dimensional max pooling with a size of 2, followed by a batch normalization layer and ReLU activation functions. Global average pooling performs on the output of the encoder. Our ResNet model consists of one convolutional residual block, including two 1-dimensional convolution layers with kernel sizes 4 and 2, with 128 and 4 filters, with strides of 4 and 2, respectively. The main block is comprised of three 1-dimensional convolution layers, each with a kernel size of 3 with 512, 128, and 4 filters. Each layer is followed by a batch normalization layer and ReLU activation functions.
In Table II, we compare our scheme in terms of the performance against Baseline 4 (same settings as Baseline 4 in Table 1 [in the main paper]), where the entire architecture is trained on the edge. Let us first consider the Transformer model. In Baseline 4, both base and head architectures are trained on the edge, i.e., a total of 128,882 parameters. This results in an accuracy and an F1-score of $83.4\%$ and $82.3\%$, respectively. When using our approach, however, we reduce the number of parameters trained on the edge to 8,234, which is roughly $6.4\%$ of the entire architecture, while obtaining a similar performance. In the case of ResNet model, Baseline 4 trains the entire architecture on the edge, which has a total of 145,226 parameters. This results in an accuracy and an F1-score of $80.9\%$ and $81.0\%$, respectively. When using our approach, however, we reduce the number of parameters trained on the edge to 8,234, which is roughly $5.7\%$ of the entire architecture, while obtaining on par performance (less than $1\%$ difference). This evaluation across diverse architectures underscores the adaptability and effectiveness of our proposed approach, showcasing its applicability beyond a specific network configuration considered earlier.
TABLE II: Evaluation of the impact of network architectures (average of 10 runs)
@@
(a) Scenario one
(b) Scenario two
(c) Scenario three
IV-B2 Performance and Overhead
In this section, we evaluate the overheads of our proposed edge FL framework.
Let us first look into the impact of knowledge distillation for several student network architectures. The results are shown in Table III. Each number represents the average of 10 runs. The results show the trade-off between the size of the student network architecture, which is captured by the number of parameters, and the prediction performance metrics, captured by the precision, recall, accuracy, and F1-score. We have selected Student 2 as a favorable trade-off between the network size and prediction performance and reported the results in Table I, which will be discussed further in this section.
As described in Section IV-A3, the base and head architectures in the Student 2 network have three layers (with 135,052 trainable parameters) and one layer (with 8,234 trainable parameters), respectively. Our results show that, it is possible to reach a prediction performance of $81.2\%$ by training less than $6\%$ of the parameters in the entire deep neural network on the edge IoT device, compared to training the entire deep neural network with $82.7\%$. Therefore, our proposed framework enables efficient training of neural networks on resource-constrained IoT devices, without any major loss in terms of prediction performance.
TABLE III: Evaluation of the impact of knowledge distillation with several student network architectures (average of 10 runs)

### v2

@@ Abstract
keywords:
[…]
1 Introduction
Over the past few decades, we have been witnessing a dramatic increase in healthcare costs. In the United States of America, for instance, the healthcare costs have risen from 27.2 billion dollars (only 5 percent of GDP) in 1960 to 41.2 billion dollars (19.7 percent of GDP) in 2020 [1, 2]. Current practice in the healthcare system places a heavy burden on the society, in terms of healthcare expenditures, due to its treatment-oriented philosophy. Mobile-health and wearable technologies offer a promising solution to pervasive healthcare by removing the constraints with respect to time and location [3]. Published in Future Generation Computer Systems (https://doi.org/10.1016/j.future.2024.07.035)
@@ 1 Introduction
The majority of the state-of-the-art detection techniques for such adverse health events, i.e., epileptic seizures, strokes, heart attacks, adopt modern ML algorithms, e.g., deep neural networks [10, 11, 12, 13]. However, training reliable deep neural network models requires a large amount of data, i.e., the so-called Big Data. As a result, the training is often performed based on the data collected from many patients, which are stored and distributed on patients’ mobile devices and wearable sensors. Traditionally, all training data samples are assumed to be present where the ML algorithm runs, i.e., the assumption was that the learning algorithm has access to all training data samples throughout the entire learning process. Therefore, to learn a model by adopting conventional solutions, each party holding a portion of the data is required to transfer all its training data in raw format to a center for the learning process. However, such an approach raises major privacy concerns in connection to sharing sensitive personal/patient data [14, 15].
@@
The paper is structured as follows. Section 2 reviews the state of the art in FL, including in the context of the IoT systems and health applications, with a main focus on privacy. Section 3 illustrates the proposed framework for privacy-preserving edge FL, designed for resource-constrained mobile-health and wearable technologies. Section 4 evaluates the proposed framework on real-world IoT applications, including implementation of our proposed framework on Amazon’s AWS cloud. Finally, Section 5 serves as the conclusion of this article.
2 State of the Art
@@
In the past few years, a large number of studies consider FL to address their concerns related to privacy and overhead, particularly in the healthcare domain [39, 40, 12]. FL methods are usually based on neural network algorithms, but the term has also been used for other ML algorithms recently, e.g., for tree-based algorithms [41, 42], for algorithms that are both gradient-based and tree-based [43, 44, 45, 46, 47], or for other algorithms [48, 49]. Several literature reviews of the state-of-the-art FL techniques are available in [50, 51, 52].
@@
A comprehensive study on the security and privacy of FL is performed in [86]. The authors extensively discuss the vulnerabilities and threats in FL. For instance, one concern in FL is data poisoning [87], which is relevant when the attacker incorporates malicious data points to training data in order to maximize the error. Membership inference attack [80, 81] is another concern that enables determining whether a specific record was included in the data used to train a model.
@@
3 Privacy-Preserving FL
In this section, we describe our proposed privacy-preserving FL framework for mobile-health and wearable technologies. Section 3.1 will briefly discuss the privacy model we consider in this paper. Then, in Section 3.2, we discuss our proposed privacy-preserving edge FL framework for mobile-health and wearable technologies.
3.1 Privacy Model
@@
3.2 Privacy-Preserving Edge FL
We shall now present our proposed privacy-preserving edge FL in two steps. In the first step, in Section 3.2.1, we illustrate the initialization procedure among healthcare centers to train an initial base neural network model in a privacy-preserving fashion. Then, in Section 3.2.2, we discuss our privacy-preserving FL framework for mobile-health and wearable technologies, which reduces the complexity by exploiting the the initial base neural network model trained in Section 3.2.1.
3.2.1 FL Base Model
In the initialization phase, we assume $N$ hospitals or healthcare centers are involved in training an initial generic model for epileptic seizure detection. The training data is distributed among all healthcare centers involved and the training is required to be performed in a secure and distributed fashion, since the personal/patient medical data is generally considered sensitive and may not leave hospital premises due to regulatory restrictions. This scenario is shown in Figure 1.
@@ 3.2.1 FL Base Model
Given a deep neural network model and the data, the goal is to train a robust and reliable model based on the training data available on all healthcare centers. The model (hypothesis) is shown by $h_{\theta}(X)$, in which $\theta$ is the set of the network’s parameters that are optimized during the process of learning, and $X$ is the input to the network. The objective is to minimize the loss represented by the following function:
@@
where $m$ represents the number of training data samples. $X^{(i)}$ denotes the attributes for the $i^{\textrm{th}}$ training data sample, and $y^{(i)}$ is the target value for the $i^{\textrm{th}}$ training data sample. Having the loss function $J(\theta)$, the model parameters $\theta$, the training data $X$, and the labels $y$, we compute the gradients for updating the model parameters. Having the gradients, we update the model parameters based on an optimization algorithm, e.g., the Adam optimizer. This process summarizes the conventional centralized training procedure for DNNs.
@@
The standard FL scheme, however, while reducing the privacy concerns, does not guarantee privacy. Therefore, the data-holder parties may not directly share their contribution, partial gradients, with other parties for privacy and legal concerns [76, 77, 78, 79]. Several studies incorporate noise to data-holder parties’ contribution in order to provide differential privacy [84]. However, this degrades the prediction performance as there is a trade-off between privacy and data utility in differential privacy [85]. In contrast, here, the optimizer merely requires the global gradient to update the network parameters and we do not require parties to share their partial gradients in their raw format. Therefore, in this study, for calculating the global gradient, we employ a secure aggregation technique. In essence, our proposed framework masks the contributions of each data-holder party to preserve privacy. Towards this, each party includes its local masks to the partial gradients and only then shares the results with the mediator (a central server orchestrating the training process):
@@ 3.2.2 FL on Edge
In the first step, similar to the previous section, the key exchange protocol via the Diffie-Hellman scheme [108] is adopted to exchange secret shares among the parties in a pair-wise fashion. The Diffie-Hellman key exchange scheme is, however, relatively complex when it comes to mobile-health and wearable technologies. To alleviate this, instead of executing the Diffie-Hellman key exchange scheme in every iteration of the learning process, we propose to share a secret random seed pair-wise among the parties in the initialization phase. This seed will then be used to generate secret masks during the entire process of privacy-preserving FL.
@@
The key challenge here is that training this model from scratch incurs major overheads for resource-constrained mobile-health and wearable technologies. On the one hand, while the initial generic neural network model trained collaboratively by the healthcare centers (in Section 3.2.1) is tailored to the high-quality data from the high-tech hospital equipment, this model can still be used as a basis for the data acquired by the mobile-health and wearable technologies. This will allow faster convergence of the model, hence reducing the communication and computation overheads of the training.
Unfortunately, training the entire neural network, even starting from the already trained models in the initial step, is still computationally infeasible for the resource-constrained mobile-health and wearable technologies. To address this challenge, inspired by the transfer learning techniques, we propose to only refine the model partially, where only the last few layers are fine-tuned to the data acquired by the mobile-health and wearable devices. More specifically, the initial generic model, denoted by $M$, is already available. As shown in Figure 2, the model $M=[B\rightarrow H]$ consists of two parts: the base architecture, denoted by $B$, and the head architecture, denoted by $H$. To reduce the overheads of training on resource-constrained mobile-health and wearable devices, we only train the head architecture $H$ on the mobile-health devices in a federated and privacy-preserving fashion. The base architecture $B$, while trained on high-quality data from high-tech hospital equipment, is still able to capture patterns which are also relevant in the case of the data acquired by the mobile-health and wearable devices. Mathematically, this is captured by the following equation:
@@
Our proposed framework also offers the possibility of personalization, to enable the realization of the precision medicine paradigm. Once the FL process discussed above is completed, the head $H$ can be immediately refined based on the data locally available to each patient. In this way, the parameters in the head $H$ architecture will be fine-tuned to the specific profile of each patient, enabling precision detection of health pathology, as we show experimentally in the next section.
In summary, our proposed privacy-preserving edge FL framework is specifically designed for resource-constrained IoT, mobile-health, and wearable technologies, not only from the FL perspective, but also from a privacy perspective. In the next section, we will evaluate our proposed privacy-preserving edge FL framework with respect to prediction performance, privacy, and overheads.
4 Experimental Setup and Results
This section presents the experimental results and evaluation of our proposed framework with respect to classification performance, overhead, and privacy, which are the main criteria for the assessment of privacy-preserving data mining algorithms [110]. In Section 4.1, we discuss the experimental setup that includes the specification of the data in our experiments and the architecture of our neural network. In Section 4.2, we evaluate our proposed privacy-preserving edge FL framework in terms of overheads, complexity, latency, and privacy, based on an implementation of the proposed framework on Amazon’s AWS cloud platform.
4.1 Experimental Setup
4.1.1 Epilepsy Dataset
In this work, we consider the CHB-MIT database [21, 22] that contains Electroencephalogram (EEG) signals from 23 epilepsy patients with intractable seizures. All recordings are collected from children and young adults in the 1.5–22 age range. For collecting data, the international 10-20 system is used, which exploits EEG electrodes with a frequency of 256 Hz and 16-bit resolution. The dataset is annotated by the medical experts and contains a total of 182 seizures. These EEG signals are sampled at $F_{s}=256$ Hz, with 16-bit resolution. Here, we consider only the two channels $T7F7$ and $T8F8$, as used in e-Glass wearable system [7], which have been shown to be important for the detection of epileptic seizures.
4.1.2 Seizure Detection Performance Metrics
@@
where $FP$, $TN$, $TP$, and $FN$ capture the number of false positives, true negatives, true positives, and false negatives, respectively.
4.1.3 Deep Neural Network’s Architecture
As described in Section 3, we have two phases of FL. In the first phase, since the training is performed on decentralized data stored in several hospitals, the assumption is that we do not have limitations in regard to the resources required for training our model. Therefore, we can have more trainable parameters in the network’s base, if increasing the size of the base improves the classification performance. On the other hand, in the second phase, since the training is performed on resource-constrained devices, we desire to have a lower number of network parameters to train. To this end, we designed our network to have three convolution layers with 135,052 trainable parameters in the base to more effectively extract the features from input EEG signals. For the head of our network, we merely employed one dense layer with 8,234 trainable parameters.
Figure 3 shows the architecture of the base and head of the neural network we used in our experiments. [^1] The network architecture was plotted using [111]. We designed the base to have three layers of one-dimensional convolution, each followed by batch normalization, activation function, and dropout. The convolution kernel size and stride are set to 2, and the number of filters for the first, second, and third layers are set to 512, 128, and 4, respectively. The activation functions in the base are rectified linear unit (ReLu) functions, and the dropout rates are set to 30%. For the head of the network, we consider a dense layer with eight units, followed by batch normalization, ReLu activation function, and dropout. Then, we have an output layer that contains two units with softmax activation functions.
@@ 4.1.3 Deep Neural Network’s Architecture
Table 1: Quantitative comparison of the proposed framework against the baseline techniques (average of 100 runs)
@@
4.2 Experimental Results
4.2.1 Prediction Performance
In this section, we evaluate our proposed edge FL framework in terms of prediction performance and for the case of epileptic seizure detection. We perform the experiments 100 times and report the average value in Table 1. Consistent with the CHB-MIT dataset, we consider 23 patients and each with his/her own data locally and the e-Glass system [7]. We consider four baseline approaches as follows:
@@ 4.2.1 Prediction Performance
Next, our proposed framework also provides the possibility of personalization. In this setting, we follow the same exact procedure as discussed in our proposed framework. After the FL is finalized, we fine-tune the head architecture with the local data of the patient to develop a personalized model for that patient. We have repeated this experiment for all patients to obtain a personalized model for each patient. The last row of Table 1 shows the average results of the personalization for all patients. The results indicate that the personalized models outperform their generic counterpart by $7.0\%$ and $6.2\%$ in terms of accuracy and F1-score, respectively. This demonstrates that our proposed framework supports personalization to significantly improve the prediction performance.
Table 2: Evaluation of the impact of network architectures (average of 10 runs)
@@
Finally, here, we show that the proposed approach is not limited to the architecture considered in Figure 3. To demonstrate this, we consider two state-of-the-art architectures, namely Transformer [112] and ResNet [113]. Our Transformer model consists of one encoder block, including multi-head attention with 8 heads and an embedding size of 32. The feedforward part consists of two 1-dimensional convolutional layers with kernel sizes of 1, with 8 and 16 filters, and strides of 2. Each layer consists of 1-dimensional max pooling with a size of 2, followed by a batch normalization layer and ReLU activation functions. Global average pooling is performed on the output of the encoder. Our ResNet model consists of one convolutional residual block, including two 1-dimensional convolutional layers with kernel sizes 4 and 2, with 128 and 4 filters, with strides of 4 and 2, respectively. The main block is comprised of three 1-dimensional convolutional layers, each with a kernel size of 3 with 512, 128, and 4 filters. Each layer is followed by a batch normalization layer and ReLU activation functions.
In Table 2, we compare our scheme in terms of the performance against Baseline 4 (same settings as Baseline 4 in Table 1), where the entire architecture is trained on the edge. Let us first consider the Transformer model. In Baseline 4, both base and head architectures are trained on the edge, i.e., a total of 128,882 parameters. This results in an accuracy and an F1-score of $83.4\%$ and $82.3\%$, respectively. When using our approach, however, we reduce the number of parameters trained on the edge to 8,234, which is roughly $6.4\%$ of the entire architecture, while obtaining a similar performance. In the case of ResNet model, Baseline 4 trains the entire architecture on the edge, which has a total of 145,226 parameters. This results in an accuracy and an F1-score of $80.9\%$ and $81.0\%$, respectively. When using our approach, however, we reduce the number of parameters trained on the edge to 8,234, which is roughly $5.7\%$ of the entire architecture, while obtaining on par performance (less than $1\%$ difference). This evaluation across diverse architectures underscores the adaptability and effectiveness of our proposed approach, showcasing its applicability beyond a specific network configuration considered earlier.
Table 3: Evaluation of the impact of knowledge distillation with several student network architectures (average of 10 runs)
@@
(a) Scenario one
(b) Scenario two
(c) Scenario three
4.2.2 Performance and Overhead
In this section, we evaluate the overhead of our proposed edge FL framework.
Let us first look into the impact of knowledge distillation for several student network architectures. The results are shown in Table 3. Each number represents the average of 10 runs. The results show the trade-off between the size of the student network architecture, which is captured by the number of parameters, and the prediction performance metrics, captured by the precision, recall, accuracy, and F1-score. We have selected Student 2 as a favorable trade-off between the network size and prediction performance and reported the results in Table 1, which will be discussed further in this section.
As described in Section 4.1.3, the base and head architectures in the Student 2 network have three layers (with 135,052 trainable parameters) and one layer (with 8,234 trainable parameters), respectively. Our results show that, it is possible to reach a prediction performance of $81.2\%$ by training less than $6\%$ of the parameters in the entire deep neural network on the edge IoT device, compared to training the entire deep neural network with $82.7\%$. Therefore, our proposed framework enables efficient training of neural networks on resource-constrained IoT devices, without any major loss in terms of prediction performance.